Dinitrogen tetroxide decomposes to form nitrogen dioxide according to the following unbalanced equation: N2O4 (g) —-> NO2 (g)
If we start with 31.7g of dinitrogen tetroxide , what is the percent of yield of the reaction if we produce 10.2g of nitrogen dioxide?
1
Expert's answer
2020-11-25T02:20:58-0500

Solution:

The balanced chemical equation:

N2O4(g) --> 2NO2(g)

According to the chemical equation: n(N2O4) = n(NO2)/2


Moles of dinitrogen tetroxide = n(N2O4) = Mass of N2O4 / Molar mass of N2O4

The molar mass of N2O4 is 92 g mol-1.

n(N2O4) = (31.7 g ) / (92 g mol-1) = 0.3446 mol


n(NO2) = n(N2O4) × 2 = 0.3446 mol × 2 = 0.6892 mol


Moles of nitrogen dioxide = n(NO2) = Mass of NO2 / Molar mass of NO2

The molar mass of NO2 is 46 g mol-1.

The theoretical mass of NO2 = n(NO2) × M(NO2) = 0.6892 mol × 46 g mol-1 = 31.7 g


The theoretical mass of NO2 = 31.7 g

The actual mass of NO2 = 10.2 g


% yield = [(actual mass) / (theoretical mass)] ×100%

% yield = (10.2 g / 31.7 g) × 100% = 32.1766% = 32.18%

% yield = 32.18%


OR


According to the law of conservation of mass, the mass of the products in a chemical reaction must equal the mass of the reactants.

Hence,

Mass of NO2 = Mass of N2O4

The theoretical mass of NO2 = Mass of N2O4 = 31.7 g

The actual mass of NO2 = 10.2 g


% yield = [(actual mass) / (theoretical mass)] ×100%

% yield = (10.2 g / 31.7 g) × 100% = 32.1766% = 32.18%

% yield = 32.18%



Answer: 32.18% is the percent of yield of the reaction.